# Integrability versus chaos in the steady state of many-body open quantum systems

### v1

@@ I Introduction
Our starting point for describing open quantum systems is the Lindblad master equation [33, 34, 35], which assumes the bath(s) to be Markovian. In recent years, the Lindblad description has been used extensively for extended many-body systems with a well-defined thermodynamic limit, so that questions like thermalization and integrability are relevant. The investigation of these systems benefits from an increasing number of known exactly solvable models, such as quadratic systems  [36, 37, 38, 39, 40, 41, 42] and models that can be solved via the Bethe ansatz [43, 44, 45, 46, 47], as well as boundary-driven systems where only the steady state is known explicitly [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65].
@@
The dynamics of a Lindblad equation leads to a steady state in the long-time limit. (There might be multiple steady states, but such cases will not be relevant for this work.) As a result, two types of integrability naturally arise, which marks a prominent difference from isolated quantum systems. On the one hand, the full system in terms of its Liouvillian may be integrable. Then the steady state of the system is known as well. On the other hand, a model can have an integrable steady state, with the Liouvillian itself being nonintegrable. We are interested in gaining a better understanding of how these types of integrability relate to different criteria associated with quantum chaos. Therefore, we analyze spin chains exhibiting these types of integrability using measures of chaos: in particular, we use complex level spacing ratios [14] for the Liouvillian spectrum, the level spacing distribution of the steady-state density matrix $\rho_{ss}$, and an extension of ETH to nonequilibrium steady states [29]. This extension of ETH associates a Hamiltonian $H_{ss}$ with the steady state ($\rho_{ss}=e^{-H_{ss}}$) and inquires whether this “steady-state Hamiltonian” $H_{ss}$ obeys ETH.
@@
We find for the studied models that integrability in all cases leads to Poisson level statistics, for the Liouvillian as well as for the steady state. Furthermore, if just the steady state is integrable but not the Liouvillian, only the (real) levels of the steady state are Poissonian while the (complex) eigenvalues of the Liouvillian exhibit level repulsion. This demonstrates that chaotic dissipative dynamics do not always relax to a chaotic steady state. We also show that the steady-state Hamiltonian $H_{ss}$ exhibits ETH scaling [66] when the steady state is nonintegrable and fails to show this scaling behavior when the steady state is integrable.
@@
In addition, we examine explicitly the structure of the steady state, by characterizing the complexity of the steady-state Hamiltonian $H_{ss}$ using the notion of operator size [67, 68, 69, 70, 71, 72]. A reasonable conjecture might be that, in the cases where the steady state is integrable, $H_{ss}$ might be a familiar Hamiltonian known from the integrability literature, or at least that it has the simple form of being composed of local few-body operators. We test this idea for a few cases with integrable steady states. We demonstrate that, for the integrable steady states we have looked at, $H_{ss}$ is not simple in this sense: its expansion in Pauli strings includes many-body operators and has no strong locality. However, the integrable steady states are in some sense “more local” than the nonintegrable cases, and the shape of the operator size distribution allows us to identify steady-state integrability in several cases.
@@ II.1 Lindblad formalism and the Liouvillian
Under the assumption that the time evolution is Markovian, the Liouvillian superoperator $\mathcal{L}$ assumes the Lindblad form [33, 34, 35]
@@
Here $D$ is the dimension of the reduced Hilbert space, the jump operators $L_{k}$ account for the coupling to the environment, and the real positive coefficients $\gamma_{k}$ give the coupling strength of the different decay channels. The steady state $\rho_{ss}$ is defined as the eigenstate of $\mathcal{L}$ with zero eigenvalue.
@@
II.2 Boundary-driven spin chain models
We consider a family of homogeneous XXZ Heisenberg spin chains of length $N$ with nearest- and next-to-nearest-neighbour interactions, whose Hamiltonian is
@@ II.2.1 Model $A$
The first model we study is given by a nearest-neighbor XXX Hamiltonian ($J=\Delta=1$, $J^{\prime}=\Delta^{\prime}=0$), which is coupled to a source on one boundary and a sink on the other, and which has no dephasing ($\gamma^{\mathrm{deph}}=0$). For this model, one can construct the steady state $\rho_{ss}$ exactly [49, 50]. However, the full spectrum of the Liouvillian is nonintegrable [14]. We break the integrability of the steady state by adding next-to-nearest-neighbor interactions in the bulk Hamiltonian, setting $J^{\prime}=1$. The nonintegrable model will be referred to as $A^{\prime}$ (see Table 1).
@@ II.2.2 Model $B$
As a second model, we consider a boundary-driven XX-spin chain with additional dephasing. Here, the driving is an arbitrary coupling to sources and sinks at both boundaries together with dephasing. This model can be mapped to the Fermi-Hubbard chain with an imaginary interaction and imaginary boundary magnetic field and is thus Bethe-ansatz integrable [43]. In contrast to model $A$, both the full Liouvillian and the steady state are integrable. In this case, we can break integrability by introducing an anisotropy $\Delta$, by adding next-to-nearest-neighbor interactions $J^{\prime}$, or both (models $B1$, $B2$, and $B3$ in Table 1, respectively).
Table 1: Parameters for the specific models introduced in §II.2. Two values of the dephasing strength $\gamma^{\text{deph}}$ are used for models of type $B$. (The rationale is explained in §III.2.) The checkmarks in the last two columns indicate whether the Liouvillian or the steady state are integrable, respectively.
@@ II.3 Symmetries
Moreover, the steady state inherits the symmetry of the Liouvillian. Indeed, since $\mathcal{L}\exp\{\mathcal{S}^{z}\}|\rho_{ss}\rangle\!\rangle=\exp\{\mathcal{S}^{z}\}\mathcal{L}|\rho_{ss}\rangle\!\rangle=0$, if $\rho_{ss}$ is the unique steady state of $\mathcal{L}$ [29] then $\exp\{\mathcal{S}^{z}\}|\rho_{ss}\rangle\!\rangle=|\rho_{ss}\rangle\!\rangle$. In terms of operators in the original Hilbert space, we have
$[\rho_{ss},S^{z}]=0.$ (10)
@@ II.4 Level statistics
To study the level statistics of the Liouvillian, we use the complex level ratios (CSR) introduced in Ref. [14]:
$z_{m}:=\frac{\lambda_{NN}-\lambda_{m}}{\lambda_{NNN}-\lambda_{m}}.$ (11)
Here, $\lambda_{NN}$ and $\lambda_{NNN}$ are the nearest- and next-to-nearest neighbors of an eigenvalue $\lambda_{m}$ of the Liouvillian. The use of complex spacing ratios instead of the complex spacings themselves avoids the need to unfold the spectrum [10]. Also, as the ratios are complex, they contain additional information about the complex spectrum of the Liouvillian. Computing the spacing ratios for all levels of a block of the Liouvillian allows us to compute their distribution in the complex plane. For independent levels characteristic of an integrable spectrum, the level spacing ratios are uniformly distributed across the unit circle (2d Poisson statistics). The levels of chaotic systems repel, which leads to a suppression of the probability of ratios with small radial and angular components [14]. The amount of level repulsion is conveniently characterized by two single-number measures, namely, the radial $\langle r\rangle$ and angular $-\langle\cos\theta\rangle$ average of the spacing ratio, where $z_{m}=r_{m}e^{i\theta_{m}}$ [14].
For the steady state, we look at the effective Hamiltonian $H_{ss}$ defined as:
$\rho_{ss}=e^{-H_{ss}}.$ (12)
As before, we will use the level statistics of $H_{ss}$ to detect chaotic or regular behavior. Since the steady state ‘Hamiltonian’ $H_{ss}$ is Hermitian, the spectrum is now real. Again to circumvent unfolding, we consider level spacing ratios of the eigenvalues of $H_{ss}$, defined as [74, 75]:
@@
where $s_{m}$ are the spacing between consecutive eigenvalues of $H_{ss}$. Their approximate distributions for independent random levels and the eigenvalues of random matrices are derived in Ref. [75] using Wigner-like surmises. The repulsion of correlated levels suppresses the probability of small $\tilde{r}$ in chaotic systems.
@@ II.5 NESS ETH
Here $\bar{O}$ and $f_{O}$ are smooth functions of the mean energy $\bar{E}=(E_{m}+E_{n})/2$ and the frequency $\omega=E_{n}-E_{m}$, with $\bar{O}(\bar{E})$ equalling the microcanonical expectation value at $\bar{E}$. $S(\bar{E})$ is the thermodynamic entropy and $R_{mn}$ is a random variable with zero mean and variance one. This hypothesis states that the diagonal matrix elements coincide with the microcanonical expectation value up to subleading fluctuations, while off-diagonal elements fluctuate around zero. Importantly, the standard deviation $\sigma$ of the fluctuations decreases with the system size in the thermodynamic limit. Away from the spectral edges, this finite-size scaling is expected to obey a power law [66]
@@
$\rho_{ss}\approx\frac{1}{Z}e^{-\beta H}.$ (16)
@@
$\rho_{ss}\propto e^{-\sum_{k}\beta_{k}h_{k}+\dots},$ (17)
where $h_{k}$ and $\beta_{k}$ are, respectively, the local energy density and inverse temperature at site $k$, and the ellipsis denotes contributions by currents of conserved quantities. Hence, the steady-state Hamiltonian $H_{ss}$ defined by Eq. (12) is expected to be local, in good approximation, and to have a similar structure as $H$. This motivated the extension of ETH to nonequilibrium steady states (NESS ETH), which states that the ansatz of Eq. (14) holds for the steady-state Hamiltonian $H_{ss}$ if the steady state is nonintegrable [29]. Although the models we examine are not in the regime of weak driving, we will check whether the ansatz (14) and finite-size scaling (15) hold for $H_{ss}$.
In the models considered, the dimensions of the steady-state sector and of the full Hilbert space scale the same way with the system size $N$. For demonstrating the presence or absence of ETH scaling, we will plot the width $\sigma$ as a function of the system size $N$, which is proportional to the logarithm of the dimension of the steady-state sector and also to the logarithm of the full Hilbert space dimension. If $\sigma$ is plotted on a logarithmic scale as a function of $N$, the $\sigma\propto D^{-1/2}$ dependence appears as a straight line in such a plot.
@@ III Chaos versus integrability in boundary-driven spin chains
We obtained the eigenvalues of the boundary-driven Liouvillians by full numerical diagonalization. The same procedure also gave the steady state for smaller system sizes. For systems with $N\geq 9$ the steady state was computed by time evolving the master equation (2) until converging to $\rho_{ss}$. Using these density matrices, we studied the level statistics and NESS ETH, and compared with RMT or Poisson statistics. We summarize the results for the level statistics in Table 2 and discuss the results for the different models in detail in the following.
@@
Figure 1: Level statistics of $\mathcal{L}$ and $H_{ss}$ and NESS ETH in models $A$ and $A^{\prime}$. (a) Complex spacing ratios of the Liouvillian $\mathcal{L}$ for the steady-state sector, with system size $N=9$. (b) Level spacing ratios of the steady-state Hamiltonian $H_{ss}$ in the zero-magnetization sector, with system size $N=12$. (c) Standard deviation $\sigma$ of matrix elements of $S^{z}_{2}$ (in the eigenbasis of $H_{ss}$) calculated in a window of width $0.5$ in the center of the spectrum as a function of the system size. (d) and (e): Expectation values of the operator $S^{z}_{2}$ in the eigenstates of $H_{ss}$ for model $A$ (d) and $A^{\prime}$ (e), plotted against the energy density $\epsilon_{n}=E_{n}/N$. The dashed line denotes the microcanonical expectation value.
III.1 XXX-chain (model $A$)
@@
For both models $A$ and $A^{\prime}$, the Liouvillian is nonintegrable. As expected, the CSR distributions significantly deviate from Poisson statistics, see Fig. 1(a). While $A^{\prime}$ agrees well with the statistics of the GinUE ensemble, $A$ lies in between fully chaotic and regular statistics. Since the Liouvillian does not have any transposition symmetry [13, 14, 21], we conjecture that the deviation from RMT is a finite-size effect.
@@
For the steady-state Hamiltonian $H_{ss}$, the level statistics and NESS ETH agree with its integrability, see Figs. 1(b)–(e). Although we only show the level spacing ratios of $H_{ss}$, we expect $H_{ss}$ and $\rho_{ss}=e^{-H_{ss}}$ to share the same level statistics in the thermodynamic limit as $e^{x}$ is a smooth function; we have checked this equivalence explicitly for our models. For model $A$, the integrable steady state exhibits Poisson level statistics. Moreover, the standard deviation of the matrix elements of the observable $S^{z}_{2}$ is not exponentially suppressed with system size, see Fig. 1(c), indicating the failure of the NESS ETH ansatz. On the other hand, in model $A^{\prime}$, the breaking of the steady-state integrability leads to GUE level statistics and we find the scaling behavior at large $N$ to be consistent with the expectation of NESS ETH, namely $\sigma_{S^{z}_{2}}\sim 2^{-N/2}$, as seen in Fig. 1(c).
@@
We can tune the degree of integrability-breaking in model $A$ by varying the strength of the next-to-nearest-neighbor interactions $J^{\prime}$, see Fig. 2(a)–(b). The tuned model then interpolates between $A$ and $A^{\prime}$. As the level statistics show, see Fig. 2(a), there is a finite-size crossover from Poisson statistics at $J^{\prime}=0$ to RMT statistics for $J^{\prime}>0$. With increasing $L$, both the fluctuations around the RMT value and the window of $J^{\prime}$ for which we observe approximate Poisson statistics decrease. The scaling of the matrix elements in the steady-state eigenbasis is consistent with this crossover, see Fig. 2(b), but exhibits a small systematic deviation from the NESS ETH prediction ($\alpha=1/2$), even for large $J^{\prime}$. This is possibly a finite-size effect, since, to determine the scaling, we make a linear fit $\sigma\sim D^{\alpha}$ to the data for different system sizes. The exponent $\alpha$ is larger for smaller $L$, approaching $-1/2$ as $L$ increases, and hence is underestimated by the fit.
@@
Figure 2: Integrability breaking when tuning the strength $J^{\prime}$ of next-to-nearest-neighbour interactions, from model $A$ to model $A^{\prime}$ and beyond. (a) Average level spacing ratios for $N=8$ and $N=12$. The dashed lines correspond to the average for Poisson and GUE level statistics. (b) Scaling of the standard deviation $\sigma$ of $S_{2}^{z}$ (blue circles) and spin current operator $J^{z}=\frac{1}{N}\sum_{j=2}^{N}\sigma_{j-1}^{x}\sigma_{j}^{y}-\sigma_{j-1}^{y}\sigma_{j}^{x}$ (red squares). We plot the exponent $\alpha$ obtained by linear fitting the dependency $\sigma\propto D^{\alpha}$—see Fig. 1(c).
III.2 XX-chain with dephasing (model $B$)

### v2

@@ I Introduction
Our starting point for describing open quantum systems is the Lindblad master equation [33, 34, 35], which assumes the bath(s) to be Markovian. In recent years, the Lindblad description has been used extensively for extended many-body systems with a well-defined thermodynamic limit, so that questions like thermalization and integrability are relevant. The investigation of these systems benefits from an increasing number of known exactly solvable models, such as quadratic systems  [36, 37, 38, 39, 40, 41, 42] and models that can be solved via the Bethe ansatz [43, 44, 45, 46, 47, 48], as well as systems where only the steady state is known explicitly [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67].
@@
The dynamics of a Lindblad equation leads to a steady state in the long-time limit. (There might be multiple steady states, but such cases will not be relevant for this work.) As a result, two types of integrability naturally arise, which marks a prominent difference from isolated quantum systems. On the one hand, the full system in terms of its Liouvillian may be integrable. Then the steady state of the system is known as well. On the other hand, a model can have an integrable steady state, with the Liouvillian itself being nonintegrable. We are interested in gaining a better understanding of how these types of integrability relate to different criteria associated with quantum chaos. Therefore, we analyze spin chains exhibiting these types of integrability using measures of chaos: in particular, we use complex level spacing ratios [14] for the Liouvillian spectrum, the level spacing distribution of the steady-state density matrix $\rho_{\mathrm{ss}}$ [68], and an extension of ETH to nonequilibrium steady states [29]. This extension of ETH associates a Hamiltonian $H_{\mathrm{ss}}$ with the steady state ($\rho_{\mathrm{ss}}=e^{-H_{\mathrm{ss}}}$) and inquires whether this “steady-state Hamiltonian” $H_{\mathrm{ss}}$ obeys ETH.
@@
We find for the studied models that full integrability in all cases leads to Poisson level statistics, for the Liouvillian as well as for the steady state. Furthermore, if just the steady state is integrable but not the Liouvillian, only the (real) levels of the steady state are Poissonian while the (complex) eigenvalues of the Liouvillian exhibit level repulsion. This demonstrates that chaotic dissipative dynamics do not always relax to a chaotic steady state. We also show that the steady-state Hamiltonian $H_{\mathrm{ss}}$ exhibits ETH scaling [69] when the steady state is nonintegrable and fails to show this scaling behavior when the steady state is integrable.
@@
In addition, we examine explicitly the structure of the steady state, by characterizing the complexity of the steady-state Hamiltonian $H_{\mathrm{ss}}$ using the notion of operator size [70, 71, 72, 73, 74, 75, 76]. A reasonable conjecture might be that, in the cases where the steady state is integrable, $H_{\mathrm{ss}}$ might be a familiar Hamiltonian known from the integrability literature, or at least that it has the simple form of being composed of local few-body operators. We test this idea for a few cases with integrable steady states. We demonstrate that, for the integrable steady states we have looked at, $H_{\mathrm{ss}}$ is not simple in this sense: its expansion in Pauli strings includes many-body operators and has no strong locality. However, the integrable steady states are in some sense “more local” than the nonintegrable cases, and the shape of the operator size distribution allows us to identify steady-state integrability in several cases.
@@ II.1 Lindblad formalism and the Liouvillian
Under the assumption that the time evolution is Markovian, trace preserving and completely positive, the Liouvillian superoperator $\mathcal{L}$ assumes the Lindblad form [33, 34, 35]
@@
Here $D$ is the dimension of the reduced Hilbert space, the jump operators $L_{k}$ account for the coupling to the environment, and the real positive coefficients $\gamma_{k}$ give the coupling strength of the different decay channels. The steady state $\rho_{\mathrm{ss}}$ is defined as the eigenstate of $\mathcal{L}$ with zero eigenvalue.
@@
II.2 Boundary-driven spin-chain models
We consider a family of homogeneous XXZ Heisenberg spin chains of length $N$ with nearest- and next-to-nearest-neighbor interactions, whose Hamiltonian is
@@ II.2.1 Model $A$
The first model we study is given by a nearest-neighbor XXX Hamiltonian ($J=\Delta=1$, $J^{\prime}=\Delta^{\prime}=0$), which is coupled to a source on one boundary and a sink on the other, and which has no dephasing ($\gamma^{\mathrm{deph}}=0$). For this model, one can construct the steady state $\rho_{\mathrm{ss}}$ exactly [50, 51]. However, the full spectrum of the Liouvillian is nonintegrable [14]. We break the integrability of the steady state by adding next-to-nearest-neighbor interactions in the bulk Hamiltonian, setting $J^{\prime}=1$. The nonintegrable model will be referred to as $A^{\prime}$ (see Table 1).
@@ II.2.2 Model $B$
As a second model, we consider a boundary-driven XX spin chain with additional dephasing. Here, the driving is an arbitrary coupling to sources and sinks at both boundaries together with dephasing. This model can be mapped to the Fermi-Hubbard chain with an imaginary interaction and imaginary boundary magnetic field and is thus Bethe-ansatz integrable [43]. In contrast to model $A$, both the full Liouvillian and the steady state are integrable. In this case, we can break integrability by introducing an anisotropy $\Delta$, by adding next-to-nearest-neighbor interactions $J^{\prime}$, or both (models $B1$, $B2$, and $B3$ in Table 1, respectively).
Table 1: Parameters for the specific models introduced in Sec. II.2. Two values of the dephasing strength $\gamma^{\text{deph}}$ are used for models of type $B$. (The rationale is explained in Sec. III.2.) The checkmarks in the last two columns indicate whether the Liouvillian or the steady state are integrable, respectively.
@@ II.3 Symmetries
Moreover, the steady state inherits the symmetry of the Liouvillian. Indeed, since $\mathcal{L}\exp\{\mathcal{S}^{z}\}|\rho_{\mathrm{ss}}\rangle\!\rangle=\exp\{\mathcal{S}^{z}\}\mathcal{L}|\rho_{\mathrm{ss}}\rangle\!\rangle=0$, if $\rho_{\mathrm{ss}}$ is the unique steady state of $\mathcal{L}$ [29] then $\exp\{\mathcal{S}^{z}\}|\rho_{\mathrm{ss}}\rangle\!\rangle=|\rho_{\mathrm{ss}}\rangle\!\rangle$. In terms of operators in the original Hilbert space, we have
$[\rho_{\mathrm{ss}},S^{z}]=0.$ (10)
@@ II.4 Level statistics
To study the level statistics of the Liouvillian, we use the complex spacing ratios (CSR) introduced in Ref. [14]:
$z_{m}:=\frac{\lambda_{\mathrm{NN}}-\lambda_{m}}{\lambda_{\mathrm{NNN}}-\lambda_{m}}.$ (11)
Here, $\lambda_{\mathrm{NN}}$ and $\lambda_{\mathrm{NNN}}$ are the nearest- and next-to-nearest neighbors of an eigenvalue $\lambda_{m}$ of the Liouvillian. The use of complex spacing ratios instead of the complex spacings themselves avoids the need to unfold the spectrum [10]. Also, as the ratios are complex, they contain additional information about the complex spectrum of the Liouvillian. Computing the spacing ratios for all levels of a block of the Liouvillian allows us to compute their distribution in the complex plane. For independent levels characteristic of an integrable spectrum, the level spacing ratios are uniformly distributed across the unit circle (2d Poisson statistics). The levels of chaotic systems repel, which leads to a suppression of the probability of ratios with small radial or angular components [14]. The amount of level repulsion is conveniently characterized by two single-number measures, namely, the radial $\langle r\rangle$ and angular $-\langle\cos\theta\rangle$ average of the spacing ratio, where $z_{m}=r_{m}e^{i\theta_{m}}$ [14].
For the steady state, we look at the effective Hamiltonian $H_{\mathrm{ss}}$ defined as:
$\rho_{\mathrm{ss}}=e^{-H_{\mathrm{ss}}}.$ (12)
As before, we will use the level statistics of $H_{\mathrm{ss}}$ to detect chaotic or regular behavior. The steady states of models A and B are integrable in the sense that they can be written down efficiently (see Ref. [50, 51, 60] for model A and Ref. [43] for model B). This is different from integrability in the usual Hamiltonian sense; nevertheless, our expectation is that if a steady state is deemed “integrable” then one condition it should satisfy is to display Poisson statistics [68]. Since the steady state ‘Hamiltonian’ $H_{\mathrm{ss}}$ is Hermitian, the spectrum is now real. Again to circumvent unfolding, we consider level spacing ratios of the eigenvalues of $H_{\mathrm{ss}}$, defined as [78, 79]:
@@
where $s_{m}$ are the spacing between consecutive eigenvalues of $H_{\mathrm{ss}}$. Their approximate distributions for independent random levels and the eigenvalues of random matrices are derived in Ref. [79] using Wigner-like surmises. The repulsion of correlated levels suppresses the probability of small $\tilde{r}$ in chaotic systems.
@@ II.5 NESS ETH
Here $\bar{O}$ and $f_{O}$ are smooth functions of the mean energy $\bar{E}=(E_{m}+E_{n})/2$ and the frequency $\omega=E_{n}-E_{m}$, with $\bar{O}(\bar{E})$ equaling the microcanonical expectation value at $\bar{E}$. $S(\bar{E})$ is the thermodynamic entropy and $R_{mn}$ is a random variable with zero mean and variance one. This hypothesis states that the diagonal matrix elements coincide with the microcanonical expectation value up to subleading fluctuations, while off-diagonal elements fluctuate around zero. Importantly, the standard deviation $\sigma$ of the fluctuations decreases with the system size in the thermodynamic limit. Away from the spectral edges, this finite-size scaling is expected to obey a power law [69]
@@
$\rho_{\mathrm{ss}}\approx\frac{1}{Z}e^{-\beta H}.$ (16)
@@
$\rho_{\mathrm{ss}}\propto e^{-\sum_{k}\beta_{k}h_{k}+\dots},$ (17)
where $h_{k}$ and $\beta_{k}$ are, respectively, the local energy density and inverse temperature at site $k$, and the ellipsis denotes contributions by currents of conserved quantities. Hence, the steady-state Hamiltonian $H_{\mathrm{ss}}$ defined by Eq. (12) is expected to be local, in good approximation, and to have a similar structure as $H$. This motivated the extension of ETH to nonequilibrium steady states (NESS ETH), which states that the ansatz of Eq. (14) holds for the steady-state Hamiltonian $H_{\mathrm{ss}}$ if the steady state is nonintegrable [29]. Although the models we examine are not in the regime of weak driving, we will check whether the ansatz (14) and finite-size scaling (15) hold for $H_{\mathrm{ss}}$. This approach should not be confused with a different approach to ETH in Lindbladian systems that is currently under investigation [30, 102], which studies the statistical properties of the overlap of observables with the right eigenoperators of the Lindbladian.
In the models considered, the dimensions of the steady-state sector and of the full Hilbert space scale the same way with the system size $N$. For demonstrating the presence or absence of ETH scaling, we will plot the width $\sigma$ as a function of the system size $N$, which is proportional to the logarithm of the dimension of the steady-state sector and also to the logarithm of the full Hilbert space dimension. If $\sigma$ is plotted on a logarithmic scale as a function of $N$, then the $\sigma\propto D^{-1/2}$ dependence appears as a straight line in such a plot.
@@ III Chaos versus integrability in boundary-driven spin chains
We obtained the eigenvalues of the boundary-driven Liouvillians by full numerical diagonalization. The same procedure also gave the steady state for smaller system sizes. For systems with $N\geq 9$ the steady state was computed by time evolving the master equation (2) until converging to $\rho_{\mathrm{ss}}$. In practice this meant iterating the equation for small time steps until $\frac{||\mathcal{L}\rho||}{||\rho||}<10^{-8}$, where |||||$\cdot$||| | ⋅ | | is the Hilbert-Schmidt norm on $\mathcal{H}$. To perform the computations 80GB RAM were sufficient.
Using the obtained density matrices, we studied the level statistics and NESS ETH, and we compared them with RMT or Poisson statistics. We summarize the results for the level statistics in Table 2 and discuss the results for the different models in detail in the following.
@@
Figure 1: Level statistics of $\mathcal{L}$ and $H_{\mathrm{ss}}$ and NESS ETH in models $A$ and $A^{\prime}$. (a) Complex spacing ratios of the Liouvillian $\mathcal{L}$ for the steady-state sector, with system size $N=9$. (b) Level spacing ratios of the steady-state Hamiltonian $H_{\mathrm{ss}}$ in the zero-magnetization sector, with system size $N=12$. (c) Standard deviation $\sigma$ of matrix elements of $S^{z}_{2}$ (in the eigenbasis of $H_{\mathrm{ss}}$) calculated in a window of width $0.5$ in the center of the spectrum as a function of the system size. (d) and (e): Expectation values of the operator $S^{z}_{2}$ in the eigenstates of $H_{\mathrm{ss}}$ for model $A$ (d) and $A^{\prime}$ (e), plotted against the energy density $\epsilon_{n}=E_{n}/N$. The dashed line denotes the microcanonical expectation value.
III.1 XXX chain (model $A$)
@@
For both models $A$ and $A^{\prime}$, the Liouvillian is nonintegrable. As expected, the CSR distributions significantly deviate from Poisson statistics; see Fig. 1(a). While $A^{\prime}$ agrees well with the statistics of the GinUE ensemble, $A$ lies in between fully chaotic and regular statistics. Since the Liouvillian does not have any transposition symmetry [13, 14, 21], we conjecture that the deviation from RMT is a finite-size effect.
@@
For the steady-state Hamiltonian $H_{\mathrm{ss}}$, the level statistics and NESS ETH agree with its integrability; see Figs. 1(b)–(e). Although we only show the level spacing ratios of $H_{\mathrm{ss}}$, we expect $H_{\mathrm{ss}}$ and $\rho_{\mathrm{ss}}=e^{-H_{\mathrm{ss}}}$ to share the same level statistics in the thermodynamic limit as $e^{x}$ is a smooth function; we have checked this equivalence explicitly for our models. For model $A$, the integrable steady state exhibits Poisson level statistics. Moreover, the standard deviation of the matrix elements of the observable $S^{z}_{2}$ is not exponentially suppressed with system size; see Fig. 1(c), indicating the failure of the NESS ETH ansatz. However, in model $A^{\prime}$, the breaking of the steady-state integrability leads to GUE level statistics and we find the scaling behavior at large $N$ to be consistent with the expectation of NESS ETH, namely $\sigma_{S^{z}_{2}}\sim 2^{-N/2}$, as seen in Fig. 1(c). The outlier at $N=7$ is a result of the small system size, as there are too few eigenstates within the considered energy window.
@@
We can tune the degree of integrability-breaking in model $A$ by varying the strength of the next-to-nearest-neighbor interactions $J^{\prime}$; see Fig. 2(a)–(b). The tuned model then interpolates between $A$ and $A^{\prime}$. As the level statistics show, see Fig. 2(a), there is a finite-size crossover from Poisson statistics at $J^{\prime}=0$ to RMT statistics for $J^{\prime}>0$. With increasing $L$, both the fluctuations around the RMT value and the window of $J^{\prime}$ for which we observe approximate Poisson statistics decrease. The scaling of the matrix elements in the steady-state eigenbasis is consistent with this crossover, see Fig. 2(b), but exhibits a small systematic deviation from the NESS ETH prediction ($\alpha=1/2$), even for large $J^{\prime}$. This is possibly a finite-size effect, since, to determine the scaling, we make a linear fit $\sigma\sim D^{\alpha}$ to the data for different system sizes. The exponent $\alpha$ is larger for smaller $L$, approaching $-1/2$ as $L$ increases, and hence is underestimated by the fit.
@@
Figure 2: Integrability breaking when tuning the strength $J^{\prime}$ of next-to-nearest-neighbor interactions, from model $A$ to model $A^{\prime}$ and beyond. (a) Average level spacing ratios for $N=8$ and $N=12$. The dashed lines correspond to the average for Poisson and GUE level statistics. (b) Scaling of the standard deviation $\sigma$ of $S_{2}^{z}$ (blue circles) and spin current operator $J^{z}=\frac{1}{N}\sum_{j=2}^{N}\sigma_{j-1}^{x}\sigma_{j}^{y}-\sigma_{j-1}^{y}\sigma_{j}^{x}$ (red squares). We plot the exponent $\alpha$ obtained by linear fitting the dependency $\sigma\propto D^{\alpha}$; see Fig. 1(c).
III.2 XX chain with dephasing (model $B$)